OpenStudy (anonymous):

need help... if A + B + C = π/2 radians show that tan A tan B + tan B tan C + tan C tan A = 1 thanks!

OpenStudy (anonymous):

\[\text{Because }C=\frac\pi2-(A+B)\] Then : \[\tan C=\frac1{\tan(A+B)}\] So : \[\tan A\tan B+\tan A\tan C+\tan B\tan C= \\~~~~\tan A\tan B+\tan C(\tan A+\tan B)= \\~~~~\tan A\tan B+(\tan A+\tan B)\times\frac1{\tan(A+B)} \\~~~~\tan A\tan B+(\tan A+\tan B)\times\frac{1-\tan A\tan B}{\tan A+\tan B} \\~~~~\tan A\tan B+1-\tan A\tan B=1\]

OpenStudy (anonymous):

Smart! thank you so much!!

OpenStudy (anonymous):

You are welcome !